OpenStudy (anonymous):

Write the equation in slope-intercept form of the line that has a slope of 3 and contains the point (4, 9).

OpenStudy (anonymous):

Slope Intercept Form is: y=mx+b y=3x+b 9=3(4)+b b=-3 y=3x-3